OpenStudy (swag):

Make sure you read the directions for each part carefully. This system of equations will be used in all three parts of this question. Part 1: Solve the system by graphing. Draw the graphs on paper and indicate their intersection point. Type the ordered pair solution here as your answer to Part 1. Part 2: Solve the system using the substitution method. Show all work here and indicate the solution for the system as an ordered pair. Part 3: Solve the system using the addition method. Show all work here and indicate the solution for the system as an ordered pair.

OpenStudy (swag):

EQUATION: 4x+6y=0 and x-2y=14

OpenStudy (swag):

@victoriaaa

OpenStudy (anonymous):

is this your question or are you asking it for victoriaa

OpenStudy (anonymous):

what up tomo

OpenStudy (swag):

This is my question.

OpenStudy (swag):

Please dont chat on my question -.- @50_cent

OpenStudy (swag):

I need the explanations and process. Not just the awnser @carlaelaine

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

for the first part you need to get the two equations into y = mx+b formula first.

OpenStudy (anonymous):

solve both equations for y first @SWAG

OpenStudy (swag):

I LOVE YOU <3

OpenStudy (anonymous):

hahah :) welcome